# End-to-end transfer learning for speaker-independent cross-language and cross-corpus speech emotion recognition

### v2

@@ 1 Introduction
The emotions we daily experience determine for a large part our mental flourishing and suffering. Happiness, or psychological well-being, relies greatly on how people experience positive and negative emotions in their lives [1]. Modern Human Computer Interaction (HCI) systems use image/video, speech, and physiological signals to determine people’s emotion [2]. Vocal expression is a direct and affectionate way of expressing emotions that has the advantage of being more accessible than image/video and physiological signals, for which a careful camera positioning or a well-worn wearable device is needed. As a result, SER is widely used in many applications, such as, an in-car board system that can provide aids or resolve errors in the communication according to the driver’s emotion [3], a diagnostic tool that uses the user’s speech emotion to provide diagnostic information to the physiotherapist [4], and an assistant robot that can provide emotional communication [5].
@@
SER using data-driven models has been successful in recognising emotions [6, 7, 8, 9, 10, 11, 12]. However, many data-driven models rely strongly on the mechanism underlying the generation of the data (i.e. the independent and identically distributed (i.i.d) assumption), and may hence fail when the testing data is taken from a different distribution than the training data. Alike other speech-related tasks, SER becomes highly challenging in cross-language, cross-corpus, and cross-speaker scenarios [13, 14]. To facilitate these challenges, traditional SER approaches use low-level-descriptor features that have been selected and grouped in the Geneva Minimalistic Acoustic Parameter Set (GeMAPS) feature set, including various acoustic features such as frequency-related (e.g. pitch, formant), energy-related (e.g. loudness) and spectral (e.g. spectral slope) features. In [8], the authors use low-level-descriptor features in cross-language SER, showing that people speaking different languages may express emotion in a similar way at the low-level signal level. To extract features that contain higher-level information, the i-vector proposed in [15] for speaker verification, has been further extended to SER where the “emotion i-vector” shows robustness in an English-German cross-language SER setting [16].
@@
In recent work, the complicated feature design process is taken over by a unified DNN model that enables end-to-end learning from raw data or shallow features [9, 10, 11, 12]. Robust features are then learned from data to minimise an overall loss towards emotion recognition. Many works achieve good SER performance on a single dataset [9, 10, 11, 12], but the model performance degrades dramatically in cross-language and cross-corpus scenarios [14]. To alleviate this problem, several transfer learning techniques have been applied. In [17, 18], the authors pre-train a neural network with emotional datasets from one or two languages, then fine-tune the model with a small portion of the target language data, which shows a performance improvement compared to when the model is trained with one language and tested on another languages (i.e. in the cross-language setting). In [19], a similar idea to pre-train a DNN model using multiple emotional datasets is proposed, and the authors propose to fine-tune only a task-specific parallel residual adapter which achieves increased SER performance on each target task while keeping the amount of parameters to be updated low.
@@
Other than to adapt the pre-trained model to the target language, one can also learn language-invariant or corpus-invariant features. This can be done by transforming the input space to a common subspace among each corpus. In [20], Song firstly proposed to transform the source and target input space to a common subspace where the input source-target neighbourhood relations are preserved. Zhang and Song later updated this framework to include sparsity and add discriminative power to the learning of the subspace [21]. To enable non-linear modelling with the DNN, the transformation to the common subspace is merged into an encoder neural network with a discriminator, and the learning is adversarial between the encoder and the discriminator where the encoder transforms source and target inputs to embeddings that will fool the discriminator, and then the discriminator needs to find the true related inputs (i.e. from the source dataset or the target dataset) [22, 23]. However, these methods require the definition of the “source” and “target” language beforehand, and their application is limited to those two languages. In [24], instead of using a binary discriminator, a Wasserstein distance [25] is used to measure the distances between target and source embeddings, and the method minimises the distances between them.
@@
We propose an end-to-end transfer learning framework to facilitate cross-language SER. This method lies in-between the transfer learning method and the common subspace method. First, the proposed model uses a wav2vec 2.0 feature extractor which has been trained in a self-supervised manner on about 56,000 hours of raw speech waveforms from 53 languages. Since the pre-trained wav2vec 2.0 feature extractor learns contextual structures across various speech utterances, it may capture common speech factors among different languages, hence transforming the speech waveform inputs to a common speech subspace which is shared across languages [26], and in which the corresponding speech embeddings are obtained. Next, a statistical pooling layer is applied that pools a sequence of the embeddings into one feature vector per utterance. The utterance-level feature vectors are then reduced in dimensionality, and a Deep-WCCN operation is applied to compensate for other variabilities (i.e. other than language variabilities). Finally, the compensated feature is used to predict emotion class with a simple linear classifier (constructed by a dense neural network layer). The model is fine-tuned on three emotional speech datasets with English, German, and Chinese languages using a summation of the supervised cross-entropy loss and the original wav2vec 2.0 loss. We evaluate its leave-one-language-out performance on unseen speakers, and the experimental results show that the proposed framework largely increases the cross-language SER performance compared to two well-known feature sets extracted using openSMILE for SER. It also largely outperforms many recent approaches to both within-language and cross-language SER. Finally, additional experiments on the effectiveness of the Deep-WCCN operation, and the effects of merging small amounts of target language speech data into the training set will be presented.
@@
The paper is structured as follows. First, Section 2 provides a brief overview about the most recent studies related to our work. Then we introduce the proposed model structure in Section 3, and propose some modifications to a similar approach in Deep-WCCN [27]. In Section 4, we describe the experiment datasets and the experimental settings. After that, we will present the simulation results and discuss these in Section 5. Finally, Section 6 presents the conclusions and suggestions for future work.
@@ 2.1 SER features
Traditional speech information retrieval systems use hand-crafted features such as the Mel-Frequency Cepstral Coefficients (MFCC) features and the Linear Predictive Coding (LPC) features [28], and those features are designed to largely reduce the data dimensionality and to encode temporal and spectral structures per time frame in a speech recording. A group of low-level descriptor features such as pitch, formant frequencies and bandwidths, loudness, and spectral slope have been selected and compared in [6] for SER. Reynolds and Rose later proposed to use Gaussian Mixture Models (GMMs) to model the speakers’ MFCC distribution [29]. Each individual Gaussian component of a GMM encodes a general acoustic class (e.g., vowels, nasals, or fricatives), and the spectral shape of the acoustic class can be characterised by the GMM component mean vector and the GMM component covariance matrix. The concatenation of GMM component mean vectors is a representation that contains abstract information such as speaker identity. This high-dimensional representation (referred to as the supervector in [15]) is then used to extract the i-vector which encodes the speaker and channel information in a total variability subspace [15]. However, these hand-crafted features are not designed for SER specifically, and may contain dominant misleading factors (e.g., speaker identities, language factors) which leads to poor SER performance. This problem is tackled in [16] where a two-step approach is proposed. The approach first compensates for the speaker variability in the supvervector space, then an “emotion i-vector” that encodes the variabilities of each emotion supervector centred around a maximum likelihood emotion supervector is extracted.
@@ 2.2 End-to-end learning
In the context of SER, Trigeorgis et al. first proposed an end-to-end DNN that consists of Convolutional Neural Network (CNN) layers for feature extraction from raw audio waveforms and Long Short-term Memory (LSTM) layers to model the temporal information in the feature space [30]. In [31], the authors replace the CNN with a Time-Delay Neural Network (TDNN) to create a larger receptive field. A similar idea has been proposed in [12] where a dilated CNN having a receptive field as large as the input sequence is used for both feature extraction and temporal modelling. However, generalisability remains a problem in supervised end-to-end learning since annotated SER datasets are general of small scale and recorded in specific environments, hence a model learned from these datasets may perform poorly when the test conditions are different from the training conditions, as well as in a cross-language setting.
@@ 2.3 Self-supervised learning and transfer learning
Self-supervised learning provides an end-to-end learning pathway for constructing speech features from large quantities of data without annotations. Self-supervised learning methods first transform time-domain raw audio waveforms to an embedding space (i.e. feature space), and then aim to predict randomly masked embeddings from past embeddings or adjacent embeddings [32, 33, 34, 26]. This learning scheme forces the DNN to learn intrinsic contextual structure from the data rather than modelling noise or irrelevant factors, because on a large scale, noise will not contain useful information to predict neighbouring embeddings, thus will be suppressed.
The speech features obtained from self-supervised learning can either be used directly in relevant tasks, or can be fine-tuned in a transfer learning framework. In the latter case, a few extra layers, which map the features into predictions, are added to the trained self-supervised feature extractor. Then, the predictions are evaluated with a supervised loss, and finally the entire set of DNN model parameters is updated with a supervised training dataset. An example in SER is [35] where the authors use a concatenation of several layer outputs of the wav2vec 2.0 pre-trained model for monolingual SER. However, the authors fine-tune the model on the Automatic Speech Recognition (ASR) task using a dataset that only contains neutral speech which may not be as efficient as fine-tuning the model directly on an emotive speech dataset and on a SER task.
@@ 3 Proposed transfer learning method with Deep-WCCN for cross-language SER
The proposed method is based on the wav2vec 2.0 self-supervised learning model [34]. This model has been pre-trained and used in cross-language speech recognition and has been shown to deliver speech representations that are shared across languages [26]. We first give the details of the wav2vec 2.0 model in Section 3.1, then we will present how to use the wav2vec 2.0 model in the front-stage of an end-to-end SER system, together with a modified Deep-WCCN layer inspired by similar work in [27] (in Section 3.2) and finally how to fine-tune the model under a supervised scheme in Section 3.3.
@@
The wav2vec 2.0 model learns contextual representations from speech waveforms. The model consists of a convolutional feature extractor $f$, a transformer-based sequence model $e$ and a vector quantiser module [26]. The feature extractor $f$ first transforms the raw speech waveform $\mathcal{X}$ into a sequence of $d_{z}$-dimensional latent speech representations $\bm{Z}=[\bm{z}_{1},\bm{z}_{2},...,\bm{z}_{T}]$ for $T$ time-stamps. Then the latent speech representations on one hand are fed into the sequence modeller to build contextual representations $\bm{c}_{1},\bm{c}_{2},...,\bm{c}_{T}$, and on the other hand are fed to the vector quantiser that contains $G$ codebooks, and for each codebook, there are $V$ entries. The vector quantiser discretises the latent representations and linearly transforms them into a sequence $\bm{q}_{1},\bm{q}_{2},...,\bm{q}_{T}$. The feature extractor $f$, the sequence model $e$ and the vector quantiser module are trainable and their parameters are optimised through back-propagation with two loss functions:
@@
    $\mathcal{L}_{m}=-\mathrm{log}\frac{\mathrm{exp}(\mathrm{sim}(\bm{c}_{t},\bm{q}_{t})/\kappa)}{\sum_{\tilde{\bm{q}}\in\bm{Q}_{t}}\mathrm{exp}(\mathrm{sim}(\bm{c}_{t},\bm{\tilde{q}}_{t})/\kappa)}$ (1)
@@
    $\mathcal{L}_{d}=\frac{1}{GV}\sum_{g=1}^{G}-H(\bar{p}_{g})=\frac{1}{GV}\sum_{g=1}^{G}\sum_{v=1}^{V}\bar{p}_{g,v}\mathrm{log}(\bar{p}_{g,v})$ (2)
@@
$\mathcal{L}_{ssl}=\mathcal{L}_{m}+\alpha\mathcal{L}_{d}$ (3)
To accommodate for cross-language speech variations, we use a pre-trained wav2vec 2.0 model that has been trained on 56,000 hours of speech in 53 different languages. This model is denoted as XLSR-53 and the details can be found in [26].
@@ 3.2 Deep-WCCN
The Within-Class Covariance Normalisation (WCCN) was introduced in [36], and used in speaker recognition and verification applications [36, 15]. The WCCN approach aims to learn a feature map that minimises upper bounds on the false positive and false negative rates in a linear classifier. First, let $\bm{W}=[\bm{w}_{1},\bm{w}_{2},\dots,\bm{w}_{N}]$ denote a set of $N$ $d$-dimensional feature vectors belonging to $C$ different classes $c\in\{1,\dots,C\}$. In [15], $\bm{W}$ is a set of i-vectors and $C$ is the number of different speakers. Differently in this paper, we define $\bm{W}$ to represent a set of intermediate features of the training set extracted from the front-end DNN feature extractor, and $C$ is the total number of emotion classes.
@@
The conventional WCCN pools the necessary statistics from the feature vectors of the entire training set, which is not compatible with mini-batch training in DNNs. In [27], the authors propose to estimate $\bm{S}_{w}$ with mini-batches, and to maintain a moving average $\hat{\bm{A}}$ of the corresponding mini-batch projection matrix. We believe that this moving average operation proposed in [27], which is equivalent to a weighted average of new and old batches, where the old batches are subject to exponentially decaying weights, is not the best design choice for random mini-batch Stochastic Gradient Descent (SGD) training in DNNs. Therefore, we first define the class covariance matrix $\hat{\bm{S}}_{w,c}$, which is estimated on mini-batch, and also $\hat{\bm{S}}_{w}$, which is the average of $\hat{\bm{S}}_{w,c}$ across the classes. Then we propose to maintain a cumulative average $\bar{\bm{S}}_{w}$ for $\hat{\bm{S}}_{w}$,
$\bar{\bm{S}}_{w}=\frac{N_{tot}}{N_{tot}+1}\bar{\bm{S}}_{w}+\frac{1}{N_{tot}+1}\hat{\bm{S}}_{w}$ (6)
where $N_{tot}$ is an accumulator that counts the total number of mini-batches during training. Next, we add a spectral smoothing term to the within-class covariance estimation similar to [36]:
@@
One problem of the pre-trained XLSR-53 model is that the training sets mostly consist of neutral speech, and as a consequence the pre-trained model may capture inadequate emotion structures embedded in the speech data. Therefore, we propose to fine-tune the XLSR-53 model with emotional speech data. We combine the XLSR-53 pre-trained model and the Deep-WCCN with a few extra layers into a unified model that maps time-domain raw waveforms to emotion class predictions, and this unified model is fine-tuned in a supervised manner. The model overview is shown in Figure 1. After converting the speech raw waveform to the sequence of latent speech representations $\bm{Z}$, we calculate an utterance-level feature by pooling and concatenating the statistics of $\bm{Z}$ along the time dimension as done in [37],
@@
The resulting feature vector $\bm{u}$ is a $2d_{z}$-dimensional vector that proceeds to a fully-connected layer which reduces its dimensionality to a pre-defined hidden dimension $d_{h}=\frac{1}{4}d_{z}$ and which is then followed by a Rectified Linear Unit (ReLU) non-linear activation and a dropout layer. The output from the dropout layer is the intermediate feature vector that will be fed into the Deep-WCCN to reduce the within-class variances. After that, the Deep-WCCN output feature vectors are normalised to have unit norm, and linearly transformed into predictions for emotion classification.
The fine-tuning loss is a weighted summation of the wav2vec 2.0 loss $\mathcal{L}_{ssl}$ and the emotion classification log-softmax cross-entropy loss. Specifically, given one prediction $\bm{p}=[p_{1},p_{2},\dots,p_{C}]^{T}$ and its corresponding one-hot emotion class label $\bm{y}=[y_{1},y_{2},\dots,y_{C}]^{T}$ (where only the true class label is 1, and other labels are 0) the total loss is:
$\mathcal{L}_{tot}=-\sum_{c=1}^{C}y_{c}\mathrm{log}\frac{\mathrm{exp}(p_{c})}{\mathrm{exp}(\sum_{i=1}^{C}p_{c})}+\gamma\mathcal{L}_{ssl}$ (9)
where the hyper-parameter $\gamma$ controls the weight of $\mathcal{L}_{ssl}$.

### v3

@@ 1 Introduction
The emotions we daily experience determine for a large part our mental flourishing and suffering. Happiness, or psychological well-being, relies greatly on how people experience positive and negative emotions in their lives [Krueger2014]. Modern Human Computer Interaction (HCI) systems use image/video, speech, and physiological signals to determine people’s emotion [Saxena2020]. Vocal expression is a direct and affectionate way of expressing emotions that has the advantage of being more accessible than image/video and physiological signals, for which a careful camera positioning or a well-worn wearable device is needed. As a result, SER is widely used in many applications, such as, an in-car board system that can provide aids or resolve errors in the communication according to the driver’s emotion [Schuller2004], a diagnostic tool that uses the user’s speech emotion to provide diagnostic information to the physiotherapist [France2000], and an assistant robot that can provide emotional communication [Castillo2018].
@@
SER using data-driven models has been successful in recognising emotions [Schuller2009, ElAyadi2011, Xiao2017, Zhao2019a, Zhang2019, Tzirakis2018, Tang2021]. However, many data-driven models rely strongly on the mechanism underlying the generation of the data (i.e. the independent and identically distributed (i.i.d) assumption), and may hence fail when the testing data is taken from a different distribution than the training data. Alike other speech-related tasks, SER becomes highly challenging in cross-language, cross-corpus, and cross-speaker scenarios [Bhaykar2013, Hozjan2003]. To facilitate these challenges, traditional SER approaches use low-level-descriptor features that have been selected and grouped in the Geneva Minimalistic Acoustic Parameter Set (GeMAPS) feature set, including various acoustic features such as frequency-related (e.g. pitch, formant), energy-related (e.g. loudness) and spectral (e.g. spectral slope) features. In [Xiao2017], the authors use low-level-descriptor features in cross-language SER, showing that people speaking different languages may express emotion in a similar way at the low-level signal level. To extract features that contain higher-level information, the i-vector proposed in [Dehak2011] for speaker verification, has been further extended to SER where the “emotion i-vector” shows robustness in an English-German cross-language SER setting [Desplanques2018].
@@
In recent work, the complicated feature design process is taken over by a unified DNN model that enables end-to-end learning from raw data or shallow features [Zhao2019a, Zhang2019, Tzirakis2018, Tang2021]. Robust features are then learned from data to minimise an overall loss towards emotion recognition. Many works achieve good SER performance on a single dataset [Zhao2019a, Zhang2019, Tzirakis2018, Tang2021], but the model performance degrades dramatically in cross-language and cross-corpus scenarios [Hozjan2003]. To alleviate this problem, several transfer learning techniques have been applied. In [Neumann2018, Latif2018], the authors pre-train a neural network with emotional datasets from one or two languages, then fine-tune the model with a small portion of the target language data, which shows a performance improvement compared to when the model is trained with one language and tested on another languages (i.e. in the cross-language setting). In [Gerczuk2021], a similar idea to pre-train a DNN model using multiple emotional datasets is proposed, and the authors propose to fine-tune only a task-specific parallel residual adapter which achieves increased SER performance on each target task while keeping the amount of parameters to be updated low.
@@
Other than to adapt the pre-trained model to the target language, one can also learn language-invariant or corpus-invariant features. This can be done by transforming the input space to a common subspace among each corpus. In [Song2019], Song firstly proposed to transform the source and target input space to a common subspace where the input source-target neighbourhood relations are preserved. Zhang and Song later updated this framework to include sparsity and add discriminative power to the learning of the subspace [Zhang2020]. To enable non-linear modelling with the DNN, the transformation to the common subspace is merged into an encoder neural network with a discriminator, and the learning is adversarial between the encoder and the discriminator where the encoder transforms source and target inputs to embeddings that will fool the discriminator, and then the discriminator needs to find the true related inputs (i.e. from the source dataset or the target dataset) [Latif2019, Liang2019]. However, these methods require the definition of the “source” and “target” language beforehand, and their application is limited to those two languages. In [Gideon2019], instead of using a binary discriminator, a Wasserstein distance [Arjovsky2017] is used to measure the distances between target and source embeddings, and the method minimises the distances between them.
DNN models are data-driven, incorporating large amounts and a wide variety of training data generally enhances their generalisation capability. Methods that leverage large volumes of unlabelled or synthesised data are emerging. Pepino et. al. employ a large-scale self-supervised pre-trained speech representation model, wav2vec 2.0 [Baevski2020], to extract general speech features. These features are used for within-corpus SER, showing significant performance improvements over traditional features. However, cross-language and cross-corpus performance are not evaluated in this work [Pepino2021]. In [sharma2022multi], the author utilises a variety of emotional datasets from multiple languages to fine-tune the wav2vec 2.0 model within a multi-task learning framework. The use of multilingual emotional data improves SER performance across languages included in training, though cross-language and cross-corpus results are still missing. In [ma2024emotion2vec], Ma et. al. adopt a novel self-supervised learning method based on online distillation [grill2020bootstrap] and incorporate a wide range of emotional datasets. Their method learns robust and generalised emotion representations that can be applied across different emotion recognition tasks (e.g., song emotion recognition, emotion prediction in conversation, and sentiment analysis) and languages. To synthesise data for training SER models, the authors in [latif2022self] propose a Generative Adversarial Network (GAN) based framework to generate synthetic emotional data for fine-tuning the SER model.
@@
We propose an end-to-end transfer learning framework to facilitate cross-language SER. This method lies in-between the transfer learning method and the common subspace method. First, the proposed model uses a wav2vec 2.0 feature extractor which has been trained in a self-supervised manner on about 56,000 hours of raw speech waveforms from 53 languages. Since the pre-trained wav2vec 2.0 feature extractor learns contextual structures across various speech utterances, it may capture common speech factors among different languages, hence transforming the speech waveform inputs to a common speech subspace which is shared across languages [Conneau2020], and in which the corresponding speech embeddings are obtained. Next, a statistical pooling layer is applied that pools a sequence of the embeddings into one feature vector per utterance. The utterance-level feature vectors are then reduced in dimensionality, and a Deep-WCCN operation is applied to compensate for other variabilities (i.e. other than language variabilities). Finally, the compensated feature is used to predict emotion class with a simple linear classifier (constructed by a dense neural network layer). The model is fine-tuned on three emotional speech datasets with English, German, and Chinese languages using a summation of the supervised cross-entropy loss and the original wav2vec 2.0 loss. We evaluate its leave-one-language-out performance on unseen speakers, and the experimental results show that the proposed framework largely increases the cross-language SER performance compared to two well-known feature sets extracted using openSMILE for SER. It also largely outperforms many recent approaches to both within-language and cross-language SER. Finally, additional experiments on the effectiveness of the Deep-WCCN operation, and the effects of merging small amounts of target language speech data into the training set will be presented.
@@
The paper is structured as follows. First, Section 2 provides a brief overview about the most recent studies related to our work. Then we introduce the proposed model structure in Section 3, and propose some modifications to a similar approach in Deep-WCCN [Eghbal-zadeh2018]. In Section 4, we describe the experiment datasets and the experimental settings. After that, we will present the simulation results and discuss these in Section 5. Finally, Section 6 presents the conclusions and suggestions for future work.
@@ 2.1 SER features
Traditional speech information retrieval systems use hand-crafted features such as the Mel-Frequency Cepstral Coefficients (MFCC) features and the Linear Predictive Coding (LPC) features [Hafen], and those features are designed to largely reduce the data dimensionality and to encode temporal and spectral structures per time frame in a speech recording. A group of low-level descriptor features such as pitch, formant frequencies and bandwidths, loudness, and spectral slope have been selected and compared in [Schuller2009] for SER. Reynolds and Rose later proposed to use Gaussian Mixture Models (GMMs) to model the speakers’ MFCC distribution [Reynolds1995]. Each individual Gaussian component of a GMM encodes a general acoustic class (e.g., vowels, nasals, or fricatives), and the spectral shape of the acoustic class can be characterised by the GMM component mean vector and the GMM component covariance matrix. The concatenation of GMM component mean vectors is a representation that contains abstract information such as speaker identity. This high-dimensional representation (referred to as the supervector in [Dehak2011]) is then used to extract the i-vector which encodes the speaker and channel information in a total variability subspace [Dehak2011]. However, these hand-crafted features are not designed for SER specifically, and may contain dominant misleading factors (e.g., speaker identities, language factors) which leads to poor SER performance. This problem is tackled in [Desplanques2018] where a two-step approach is proposed. The approach first compensates for the speaker variability in the supvervector space, then an “emotion i-vector” that encodes the variabilities of each emotion supervector centred around a maximum likelihood emotion supervector is extracted.
@@ 2.2 End-to-end learning
In the context of SER, Trigeorgis et al. first proposed an end-to-end DNN that consists of Convolutional Neural Network (CNN) layers for feature extraction from raw audio waveforms and Long Short-term Memory (LSTM) layers to model the temporal information in the feature space [Trigeorgis2016]. In [Sarma2018], the authors replace the CNN with a Time-Delay Neural Network (TDNN) to create a larger receptive field. A similar idea has been proposed in [Tang2021] where a dilated CNN having a receptive field as large as the input sequence is used for both feature extraction and temporal modelling. However, generalisability remains a problem in supervised end-to-end learning since annotated SER datasets are general of small scale and recorded in specific environments, hence a model learned from these datasets may perform poorly when the test conditions are different from the training conditions, as well as in a cross-language setting.
@@ 2.3 Self-supervised learning and transfer learning
Self-supervised learning provides an end-to-end learning pathway for constructing speech features from large quantities of data without annotations. Self-supervised learning methods first transform time-domain raw audio waveforms to an embedding space (i.e. feature space), and then aim to predict randomly masked embeddings from past embeddings or adjacent embeddings [VandenOordDeepMinda, Schneider2019, Baevski2020, Conneau2020], or to predict the features generated though an augmented “target” network [grill2020bootstrap]. These learning schemes force the DNN to learn intrinsic contextual structure from the data rather than modelling noise or irrelevant factors, because on a large scale, noise will not contain useful information to predict neighbouring embeddings or augmented embeddings, thus will be suppressed.
The speech features obtained from self-supervised learning can either be used directly in relevant tasks, or can be fine-tuned in a transfer learning framework. In the latter case, a few extra layers, which map the features into predictions, are added to the trained self-supervised feature extractor. Then, the predictions are evaluated with a supervised loss, and finally the entire set of DNN model parameters is updated with a supervised training dataset. Examples in SER are presented in [Pepino2021, sharma2022multi, ma2024emotion2vec].
@@ 3 Proposed transfer learning method with Deep-WCCN for cross-language SER
The proposed method is based on the wav2vec 2.0 self-supervised learning model [Baevski2020]. This model has been pre-trained and used in cross-language speech recognition and has been shown to deliver speech representations that are shared across languages [Conneau2020]. We first give the details of the wav2vec 2.0 model in Section 3.1, then we will present how to use the wav2vec 2.0 model in the front-stage of an end-to-end SER system, together with a modified Deep-WCCN layer inspired by similar work in [Eghbal-zadeh2018] (in Section 3.2) and finally how to fine-tune the model under a supervised scheme in Section 3.3.
@@
The wav2vec 2.0 model learns contextual representations from speech waveforms. The model consists of a convolutional feature extractor $f$, a transformer-based sequence model $e$ and a vector quantiser module [Conneau2020]. The feature extractor $f$ first transforms the raw speech waveform $\mathcal{X}$ into a sequence of $d_{z}$-dimensional latent speech representations $\bm{Z}=[\bm{z}_{1},\bm{z}_{2},...,\bm{z}_{T}]$ for $T$ time-stamps. Then the latent speech representations on one hand are fed into the sequence modeller to build contextual representations $\bm{c}_{1},\bm{c}_{2},...,\bm{c}_{T}$, and on the other hand are fed to the vector quantiser that contains $G$ codebooks, and for each codebook, there are $V$ entries. The vector quantiser discretises the latent representations and linearly transforms them into a sequence $\bm{q}_{1},\bm{q}_{2},...,\bm{q}_{T}$. The feature extractor $f$, the sequence model $e$ and the vector quantiser module are trainable and their parameters are optimised through back-propagation with two loss functions:
@@
    $\mathcal{L}{\mathrm{{}_{m}}}=-\mathrm{log}\frac{\mathrm{exp}(\mathrm{sim}(\bm{c}_{t},\bm{q}_{t})/\kappa)}{\sum_{\tilde{\bm{q}}\in\bm{Q}_{t}}\mathrm{exp}(\mathrm{sim}(\bm{c}_{t},\bm{\tilde{q}}_{t})/\kappa)}$ (1)
@@
    $\mathcal{L}{\mathrm{{}_{d}}}=\frac{1}{GV}\sum_{g=1}^{G}-H(\bar{p}_{g})=\frac{1}{GV}\sum_{g=1}^{G}\sum_{v=1}^{V}\bar{p}_{g,v}\mathrm{log}(\bar{p}_{g,v})$ (2)
@@
$\mathcal{L}{\mathrm{{}_{ssl}}}=\mathcal{L}{\mathrm{{}_{m}}}+\alpha\mathcal{L}{\mathrm{{}_{d}}}$ (3)
To accommodate for cross-language speech variations, we use a pre-trained wav2vec 2.0 model that has been trained on 56,000 hours of speech in 53 different languages. This model is denoted as XLSR-53 and the details can be found in [Conneau2020].
@@ 3.2 Deep-WCCN
The Within-Class Covariance Normalisation (WCCN) was introduced in [Hatch2006], and used in speaker recognition and verification applications [Hatch2006, Dehak2011]. The WCCN approach aims to learn a feature map that minimises upper bounds on the false positive and false negative rates in a linear classifier. First, let $\bm{W}=[\bm{w}_{1},\bm{w}_{2},\dots,\bm{w}_{N}]$ denote a set of NN $d$-dimensional feature vectors belonging to $C$ different classes $c\in\{1,\dots,C\}$. In [Dehak2011], $\bm{W}$ is a set of i-vectors and $C$ is the number of different speakers. Differently in this paper, we define $\bm{W}$ to represent a set of intermediate features of the training set extracted from the front-end DNN feature extractor, and $C$ is the total number of emotion classes.
@@
The conventional WCCN pools the necessary statistics from the feature vectors of the entire training set, which is not compatible with mini-batch training in DNNs. In [Eghbal-zadeh2018], the authors propose to estimate $\bm{S}_{w}$ with mini-batches, and to maintain a moving average $\hat{\bm{A}}$ of the corresponding mini-batch projection matrix. We believe that this moving average operation proposed in [Eghbal-zadeh2018], which is equivalent to a weighted average of new and old batches, where the old batches are subject to exponentially decaying weights, is not the best design choice for random mini-batch Stochastic Gradient Descent (SGD) training in DNNs. Therefore, we first define the class covariance matrix $\hat{\bm{S}}_{w,c}$, which is estimated on mini-batch, and also $\hat{\bm{S}}_{w}$, which is the average of $\hat{\bm{S}}_{w,c}$ across the classes. Then we propose to maintain a cumulative average $\bar{\bm{S}}_{w}$ for $\hat{\bm{S}}_{w}$,
$\bar{\bm{S}}_{w}=\frac{N\mathrm{{}_{tot}}}{N\mathrm{{}_{tot}}+1}\bar{\bm{S}}_{w}+\frac{1}{N\mathrm{{}_{tot}}+1}\hat{\bm{S}}_{w}$ (6)
where $N\mathrm{{}_{tot}}$ is an accumulator that counts the total number of mini-batches during training. Next, we add a spectral smoothing term to the within-class covariance estimation similar to [Hatch2006]:
@@
One problem of the pre-trained XLSR-53 model is that the training sets mostly consist of neutral speech, and as a consequence the pre-trained model may capture inadequate emotion structures embedded in the speech data. Therefore, we propose to fine-tune the XLSR-53 model with emotional speech data. We combine the XLSR-53 pre-trained model and the Deep-WCCN with a few extra layers into a unified model that maps time-domain raw waveforms to emotion class predictions, and this unified model is fine-tuned in a supervised manner. The model overview is shown in Figure 1. After converting the speech raw waveform to the sequence of latent speech representations $\bm{Z}$, we calculate an utterance-level feature by pooling and concatenating the statistics of $\bm{Z}$ along the time dimension as done in [Li2021],
@@
The resulting feature vector $\bm{u}$ is a $2d_{z}$-dimensional vector that proceeds to a fully-connected layer which reduces its dimensionality to a pre-defined hidden dimension $d\mathrm{{}_{h}}=\frac{1}{4}d_{z}$ and which is then followed by a Rectified Linear Unit (ReLU) non-linear activation and a dropout layer. The output from the dropout layer is the intermediate feature vector that will be fed into the Deep-WCCN to reduce the within-class variances. After that, the Deep-WCCN output feature vectors are normalised to have unit norm, and linearly transformed into predictions for emotion classification.
The fine-tuning loss is a weighted summation of the wav2vec 2.0 loss $\mathcal{L}\mathrm{{}_{ssl}}$ and the emotion classification log-softmax cross-entropy loss. Specifically, given one prediction $\bm{p}=[p_{1},p_{2},\dots,p_{C}]^{T}$ and its corresponding one-hot emotion class label $\bm{y}=[y_{1},y_{2},\dots,y_{C}]^{T}$ (where only the true class label is 1, and other labels are 0) the total loss is:
$\mathcal{L}{\mathrm{{}_{tot}}}=-\sum_{c=1}^{C}y_{c}\mathrm{log}\frac{\mathrm{exp}(p_{c})}{\mathrm{exp}(\sum_{i=1}^{C}p_{c})}+\gamma\mathcal{L}{\mathrm{{}_{ssl}}}$ (9)
where the hyper-parameter $\gamma$ controls the weight of $\mathcal{L}{\mathrm{{}_{ssl}}}$.